# Large Language Models as Reliable Knowledge Bases?

### v1

@@ Abstract
The NLP community has recently shown a growing interest in leveraging Large Language Models (LLMs) for knowledge-intensive tasks, viewing LLMs as potential knowledge bases (KBs). However, the reliability and extent to which LLMs can function as KBs remain under-explored. While previous studies suggest LLMs can encode knowledge within their parameters, the amount of parametric knowledge alone is not sufficient to evaluate their effectiveness as KBs. This study defines criteria that a reliable LLM-as-KB should meet, focusing on factuality and consistency, and covering both seen and unseen knowledge. [^1] Seen knowledge refers to knowledge learned during training. Unseen knowledge is neither present in the model’s training data nor can be inferred from seen knowledge. We develop several metrics based on these criteria and use them to evaluate 26 popular LLMs, while providing a comprehensive analysis of the effects of model size, instruction tuning, and in-context learning (ICL). Our results paint a worrying picture. Even a high-performant model like GPT-3.5-turbo is not factual or consistent, and strategies like ICL and fine-tuning are unsuccessful at making LLMs better KBs.
@@ 1 Introduction
Large language models (LLMs), pretrained on extensive text corpora, have demonstrated the ability to implicitly encode various types of knowledge within their weights, without requiring human supervision. As a result, many recent studies Chuang et al. (2023); Yu et al. (2023); Dhingra et al. (2022); Sung et al. (2021); Wang et al. (2020) aim to analyze the relationship between LLMs and KBs, and even explore whether LLMs can replace KBs Sun et al. (2023); Mruthyunjaya et al. (2023); Heinzerling and Inui (2021).
However, whether current LLMs can serve as reliable KBs and how to evaluate their performance in this role remains largely unexplored. Existing studies Sun et al. (2023); Wang et al. (2021); Roberts et al. (2020) often implicitly assume that the LLM’s ability to retain knowledge is sufficient for it to function as a KB. Typically, these studies employ two methods: (1) converting knowledge graphs into natural language questions using templates and evaluating LLM ability to answer these questions, by measuring the amount of knowledge therein Petroni et al. (2019); Sun et al. (2023); and (2) pre-training LLMs on passages/triples containing knowledge and then assessing their ability to answer related questions, thereby quantifying their knowledge retention Wang et al. (2021); He et al. (2024). While these methods demonstrate that LLMs can recall knowledge, memorizing facts is not the sole criterion for being a reliable KB AlKhamissi et al. (2022).
What criteria then should a LLM meet to function as a reliable KB? Discussion on this topic has been limited, and there is no agreement on the definition of these criteria. AlKhamissi et al. (2022) argue that LLMs ought to excel at five aspects (i.e., access, edit, consistency, reasoning, explainability, and interpretability) if they are to be considered a KB. However, they do not outline specific metrics to evaluate the extent to which LLMs act as KBs. Besides, we argue that evaluating LLMs against the characteristics of KBs may not be entirely appropriate due to their different data storage structures. Instead, we should consider the specific properties of LLMs when assessing their suitability as KBs.
Our research addresses these gaps and aims to establish a more nuanced understanding of LLMs-as-KBs. We evaluate and compare the reliability of different LLMs functioning as KBs in answering factoid questions. Specifically, our work makes the following contributions:
  1. 1.
    We define how to assess the reliability of LLMs-as-KBs, and propose metrics that consider the differences between KBs and LLMs and align with the distinct nature of LLMs. We consider two aspects, factuality (i.e., the ability to provide factual responses) and consistency (i.e., the ability to provide consistent responses for questions involving the same knowledge), and cover the evaluation on both seen and unseen knowledge.
  2. 2.
    To evaluate consistency, we propose a novel method which computes the probability that a LLM can consistently provide response $r$ for question $q$.
  3. 3.
    To evaluate LLMs on unseen knowledge, we create a new QA dataset UnseenQA, with a knowledge cutoff date before April 13, 2024.
  4. 4.
    We evaluate 26 popular LLMs in their ability to function as reliable KBs, and discuss the influence of model size, instruction-tuning, and ICL on their performance.
@@
Figure 1: An example illustrating three answer types: correct, wrong, and uninformative. Focusing only on the correct rate incorrectly suggests that Model B is better, even though Model A is more reliable with a similar correct rate and a much lower wrong rate.
@@ 2 What is a Reliable LLM-as-KB?
In simple terms, a LLM functions as a reliable KB if it consistently provides factual responses. Evaluating the reliability of LLMs as KBs primarily involves assessing two critical dimensions, namely factuality and consistency.
@@ 2.1 Factuality
Factuality refers to the quality of being factual or based on fact. Run-of-the-mill KBs, stored on physical servers or cloud platforms, deliver information directly in response to queries. If the requested data is unavailable, these systems typically return a null response. In contrast, LLMs are probabilistic models that excel at next word prediction based on the given context, rather than storing explicit information in defined locations. This architecture allows LLMs to generate responses that seem plausible, regardless of whether the content was included in their training data. Consequently, LLMs typically produce three types of responses: correct, uninformative, and wrong.
Ongoing efforts Chen et al. (2023); Wang et al. (2024) to evaluate the factuality of LLMs often hinge on the models’ correct rates in responding to factual QA datasets. However, this approach has notable limitations. Firstly, many studies Lin et al. (2022); Sun et al. (2023) do not specify whether the dataset’s scope of knowledge was included in the LLM’s training data. This omission can lead to unfair comparisons between models, especially if the knowledge being tested is within the training scope of one model but not another. Secondly, the assumption that a higher correct rate indicates greater factuality is problematic. For example, consider models A and B shown in Figure 1, and assume they are being evaluated on a test dataset covering knowledge they both have seen during training. In this case, focusing solely on the correct rate might erroneously suggest that Model B is more factual, despite Model A being more reliable due to its similar correct rate and significantly lower wrong rate.
Given these issues, we propose the following criteria for the factuality of LLMs-as-KBs:
CRITERION 1.1
For seen knowledge, a factual LLM should demonstrate a high correct rate and a low wrong rate.
CRITERION 1.2
For unseen knowledge, a factual LLM should demonstrate a high uninformative rate.
We next proceed to define evaluation metrics that operationalize these criteria. Let $\boldsymbol{M}$ denote a LLM. Let $D_{\text{seen}}$ denote a QA dataset containing $N$ open-ended factoid questions pertaining to knowledge the LLM ought to have seen during training. Let $D_{\text{unseen}}$ denote a QA dataset with $L$ open-ended factoid questions covering unseen knowledge. We further assume the LLM’s response to $D_{\text{seen}}$ will be correct, uninformative, or wrong, while its response to $D_{\text{unseen}}$ will be either uninformative or wrong.
@@
Figure 2: Illustration of LLM inconsistency with davinci-002 (temperature is set to 0). Questions in the top focus on seen knowledge, with probability distribution mass concentrated on one prediction. Questions in the bottom focus on unseen knowledge, where the distribution is more even. Drawing from such a distribution inevitably leads to inconsistencies.
@@ 2.2 Consistency
Consistency refers to the quality of always behaving in the same way or having the same opinions. KBs are designed with this principle in mind. In fact, there are efficient algorithms Andersen and Pretolani (2001) which detect and resolve conflicts within KBs, thus ensuring consistent outputs for queries on the same facts.
It is well-known that LLMs often exhibit inconsistencies in their responses Elazar et al. (2021); Wang et al. (2022). Current research Elazar et al. (2021); Jang et al. (2022); Hagström et al. (2023) evaluates LLM consistency through their performance on benchmarks involving paraphrasing, negation, or multilingual variations. A model is considered superior if it responds consistently across a broader range of data samples. In this work, we argue that it may be too strict to expect LLMs to be always consistent when responding to fact-based questions. Unlike explicit KBs which store information at a fixed location, LLMs operate probabilistically. In theory, if the context has been learned during training, the probability distribution for the prediction will be concentrated; otherwise, it will be more uniform. Drawing from a uniform probability distribution inevitably leads to inconsistencies. As shown in Figure 2, even with greedy decoding, slight biases in the distribution can cause fluctuations in the selection of the top probable words.
Given their probabilistic nature, we do not expect LLMs to always behave consistently. We acknowledge that inconsistencies can cause confusion in practical applications and propose to monitor model behavior through post-processing which we argue is more realistic than expecting a probabilistic model to be perpetually consistent. We thus propose the following consistency criteria:
CRITERION 2.1
The model is expected to be consistent when it produces correct responses.
CRITERION 2.2
The model is expected to be inconsistent when it produces wrong responses.
We next define evaluation metrics corresponding to the criteria above. Let $q$ refer to a question in either $D_{\text{seen}}$ or $D_{\text{unseen}}$, and $r$ denote model $M$’s response to $q$. Inspired by Zheng et al. (2024), we measure consistency based on multiple-choice questions (MCQs). As shown in Figure 3, we employ gpt-3.5-turbo-insruct to generate a set of distractor options similar to response $r$, and then create a group of MCQs. The consistency score for data point $(q,r)$ is calculated as:
$Cons(q,r)=\frac{\sum_{i=1}^{X_{\text{MCQs}}}[R_{i}=r]}{X_{\text{MCQs}}}$ (4)
where $X_{\text{MCQs}}$ is the total number of MCQs, $R_{i}$ is model $M$’s response for the $i$-th MCQ, and the expression $[R_{i}=r]$ yields 1 when the model’s response $R_{i}$ matches its original response $r$, and 0 otherwise. The consistency score $Cons(q,r)$ ranges from 0 to 1.
@@
measures the consistency of the model when it provides correct responses. It is defined as:
@@
measures the consistency of an LLM when it provides wrong responses. It is defined as:
@@
where $C_{\textit{wrong}}^{s}$ refers to the consistency of a LLM when it provides wrong responses to questions about seen knowledge, and $C_{\textit{wrong}}^{u}$ refers to the consistency of a LLM when it provides wrong responses to questions about unseen knowledge:
@@
where $r^{(w)}$ refers to responses labeled as wrong, and $q^{(w)}$ is the corresponding question. $N_{wrong}$ and $L_{wrong}$ are counts of wrong responses to $D_{\text{seen}}$ and $D_{\text{unseen}}$, respectively. $C_{\textit{wrong}}$ ranges from 0 to 1. Based on CRITERION 2.2, a lower $C_{\textit{wrong}}$ is desirable.
Figure 3: Example computation for consistency score $Cons(q,r)$. The LLM’s original answer is shown in green, while distractors are red.
@@ 2.3 Reliability (Factuality and Consistency)
A reliable LLM-as-KB should then be assessed against factuality and consistency. Based on the criteria defined above, a LLM-as-KB is reliable if it meets the following:
CRITERION 3.1
For seen knowledge, a LLM should have a high rate of consistently correct responses and a low rate of consistently wrong responses.
CRITERION 3.2
For unseen knowledge, a LLM should have a high rate of uninformative or inconsistent responses.
We quantify these criteria with two metrics.
@@
NCCR ranges from $-1$ to $1$. A higher NCCR indicates a LLM is more reliable on seen knowledge. A negative NCCR suggests the model provides consistently wrong responses, while a positive NCCR suggests a preference for consistently correct responses.
@@
assesses whether the model is likely to provide uninformative or inconsistent wrong responses for questions about unseen knowledge. It is defined as:
@@
LLM Zero-Shot Four-Shot Four-Two
NCCR IUR AVG NCCR IUR AVG NCCR IUR AVG
gpt-3.5-turbo 34.1 95.7 62.7 35.5 99.0 66.7 32.1 99.8 65.8
llama2chat-70b 20.4 98.9 59.1 17.7 99.9 58.8 16.4 100.0 58.2
llama3-70b 30.2 71.6 36.7 33.1 69.2 35.7 33.9 98.8 65.8
Table 1: Most reliable LLMs across three prompt settings. Four-Two refers to the four-shot with two unsure shots setting. All numbers shown are percentages. AVG represents average NCCR and (normalized) IUR ($2*\text{IUR}-100$) scores (both scores use the same scale). We treat seen and unseen knowledge equally, however, in practice, NCCR and IUR can be assigned different weights.
@@ 3 Experimental Setup
3.1 Datasets
To evaluate the performance of LLMs on seen knowledge, we collated SeenQA, a composite dataset containing 3,000 questions sourced from the test sets (or development sets, where test sets were unavailable) of Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), and PopQA Mallen et al. (2023). For assessing LLM performance on unseen knowledge, we introduce UnseenQA, a new dataset designed to ensure that models with a knowledge cutoff date before April 13, 2024 do not have access to answers. UnseenQA also includes 3,000 questions, which are derived from 20 templates covering various answer types such as number, people, time, location, and others. Our methodology for creating SeenQA and UnseenQA is detailed in Appendix A.

### v2

# How Reliable are LLMs as Knowledge Bases? Re-thinking Facutality and Consistency

@@ Abstract
Large Language Models (LLMs) are increasingly explored as knowledge bases (KBs), yet current evaluation methods focus too narrowly on knowledge retention, overlooking other crucial criteria for reliable performance. In this work, we rethink the requirements for evaluating reliable LLM-as-KB usage and highlight two essential factors: factuality, ensuring accurate responses to seen and unseen knowledge [^1] Seen knowledge refers to knowledge learned during training. Unseen knowledge is neither present in the model’s training data nor can be inferred from seen knowledge., and consistency, maintaining stable answers to questions about the same knowledge. We introduce UnseenQA, a dataset designed to assess LLM performance on unseen knowledge, and propose new criteria and metrics to quantify factuality and consistency, leading to a final reliability score. Our experiments on 26 LLMs reveal several challenges regarding their use as KBs, underscoring the need for more principled and comprehensive evaluation.
How Reliable are LLMs as Knowledge Bases? Re-thinking
Facutality and Consistency
Danna Zheng1, Mirella Lapata1, Jeff Z. Pan1,2 1 School of Informatics, University of Edinburgh, UK 2 Huawei Edinburgh Research Centre, CSI, UK dzheng@ed.ac.uk, mlap@inf.ed.ac.uk, http://knowledge-representation.org/j.z.pan/
@@ 1 Introduction
Large Language Models (LLMs), pretrained on vast text corpora, have demonstrated significant capabilities in encoding knowledge without explicit supervision. The continuous release of new LLMs, evaluated on benchmarks like TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), highlights their improving ability to answer fact-based queries. This progress has fueled interest in employing LLMs as knowledge bases (KBs) for various applications and developing techniques to edit model knowledge Wang et al. (2024c, b, a) or mitigate hallucinations Zhang et al. (2024b, a); Yu et al. (2024).
However, a critical question remains underexplored: What criteria should an LLM meet to function reliably as a KB? Current research often assumes that knowledge retention alone is sufficient Sun et al. (2023); Wang et al. (2021); Roberts et al. (2020). Existing evaluations generally follow two approaches: (1) converting knowledge graphs into natural language questions and assessing how many questions the LLM answers correctly Petroni et al. (2019); Sun et al. (2023); and (2) pretraining LLMs on knowledge-rich text and measuring their accuracy on related questions Wang et al. (2021); He et al. (2024).
@@
Figure 1: (a) An example illustrating three answer types: correct, wrong, and uninformative. Focusing only on the correct rate incorrectly suggests that Model B is better, even though Model A is more reliable with a similar correct rate and a much lower wrong rate. (b) Illustration of LLM inconsistency with davinci-002 (temperature is set to 0). Questions in the top focus on seen knowledge, with probability distribution mass concentrated on one prediction. Questions in the bottom focus on unseen knowledge, where the distribution is more even. Drawing from such a distribution inevitably leads to inconsistencies. (c) Example computation for consistency score $Cons(q,r)$. The LLM’s original answer is shown in green, while distractors are red. (d) An example illustrating how to evaluate LLM-as-KB.
These methods demonstrate that LLMs can recall information, but knowledge volume alone does not guarantee reliable performance as a KB. Beyond retention, it is essential to examine how LLMs handle factual queries—specifically, whether they respond accurately to seen knowledge and avoid making claims about unseen knowledge (factuality), and whether they provide consistent answers to questions about the same facts (consistency).
Factuality refers to the quality of being factual or based on fact. KBs hosted on servers or cloud platforms, offer precise answers or null responses when data is unavailable. In contrast, LLMs rely on probabilistic next-token prediction, which can lead to plausible but incorrect answers. As a result, LLM responses are typically correct, uninformative, or wrong. Existing methods for evaluating factuality often focus on the rate of correct answers in factual QA datasets Chen et al. (2023); Wang et al. (2024d). However, as many studies Lin et al. (2022); Sun et al. (2023) fail to specify whether the dataset’s knowledge was included in the LLM’s pre-training data, it is not possible to establish whether the model is genuinely factual. Secondly, it is misleading to equate a higher correct rate with greater factuality. As illustrated in Figure 1(a), a model with a higher correct rate might still be less factual if it produces a higher rate of wrong answers compared to a model with fewer errors overall.
Consistency refers to the quality of always behaving in the same way or having the same opinions. Traditional KBs achieve consistency through algorithms Andersen and Pretolani (2001) that detect and resolve conflicts. In contrast, LLMs frequently exhibit inconsistent behavior Elazar et al. (2021); Wang et al. (2022). Current research Elazar et al. (2021); Jang et al. (2022); Hagström et al. (2023) evaluates LLM consistency using benchmarks involving paraphrasing, negation, or multilingual variations, favoring models that maintain consistent responses across diverse samples. However, we argue that expecting LLMs to always be consistent in fact-based responses is overly rigid. Unlike KBs, which store information in fixed locations, LLMs operate probabilistically. When the context has been learned during training, the probability distribution for predictions is concentrated; otherwise, it remains more uniform. Sampling from a uniform distribution naturally leads to inconsistencies. As shown in Figure 1(b), even with greedy decoding, slight distribution biases can cause variations in the top-selected words.
Given these issues, this paper seeks to define the criteria for a reliable LLM-as-KB when handling factual queries. In evaluating factuality, we consider both seen knowledge (contained within training data) and unseen knowledge, and consider the negative effects of wrong answers. To assess performance on unseen knowledge, we introduce UnseenQA, a new dataset containing knowledge unavailable to LLMs trained before April 13, 2024. For consistency, we classify the correctness of responses and propose a novel method to compute the probability that an LLM can consistently provide the same response $r$ to a question $q$.
We evaluate 26 popular LLMs and find that: 1) GPT-3.5-Turbo achieves balanced performance in both factuality and consistency. LLMs like LLaMA3-70B may achieve a higher correct rate on seen knowledge but exhibit higher wrong rates and consistency on wrong answers. 2) There is a correlation between factuality and consistency: more factual LLMs tend to be consistent in their responses, whether correct or wrong. 3) Larger LLMs perform worse on unseen knowledge and are more consistent even when providing wrong answers. 4) Fine-tuning techniques can improve performance on unseen knowledge. However, this often comes at the expense of performance on seen knowledge. 5) In-context learning (ICL) does not improve performance on seen knowledge, as it as it generally increases or decreases both correct and wrong rates simultaneously. 6) Base LLMs tend to overestimate their knowledge on numerical and temporal questions.
We hope our work will draw the community’s attention to the multifaceted challenges of using LLMs-as-KBs, ultimately inspiring further research and innovation toward more reliable, robust, and principled methodologies.
@@ 2 What is a Reliable LLM-as-KB?
In simple terms, an LLM is a reliable KB if it consistently provides factual responses. As illustrating in Figure 1 (d), evaluating the reliability of LLMs as KBs primarily involves assessing two critical dimensions, namely factuality and consistency.
@@ 2.1 Factuality
We propose the following criteria for determining the factuality of LLMs-as-KBs:
Criterion 1.1: For seen knowledge, a factual LLM should demonstrate a high correct rate and a low wrong rate. Criterion 1.2: For unseen knowledge, a factual LLM should demonstrate a high uninformative rate.
We next proceed to define evaluation metrics that operationalize these criteria. Let $\boldsymbol{M}$ denote an LLM. Let $D_{\text{seen}}$ denote a QA dataset containing $N$ open-ended factoid questions pertaining to knowledge the LLM ought to have seen during training. Let $D_{\text{unseen}}$ denote a QA dataset with $L$ open-ended factoid questions covering unseen knowledge. We further assume the LLM’s response to $D_{\text{seen}}$ will be correct, uninformative, or wrong, while its response to $D_{\text{unseen}}$ will be either uninformative or wrong.
@@ 2.2 Consistency
We propose the following consistency criteria:
Criterion 2.1: The model is expected to be consistent in correct responses. Criterion 2.2: The model is expected to be inconsistent in wrong responses.
We next define evaluation metrics corresponding to the criteria above. Let $q$ refer to a question in either $D_{\text{seen}}$ or $D_{\text{unseen}}$, and $r$ denote model $M$’s response to $q$. Inspired by Zheng et al. (2024), we measure consistency based on multiple-choice questions (MCQs). As shown in Figure 1 (c), we employ gpt-3.5-turbo-insruct to generate a set of distractor options similar to response $r$, and then create a group of MCQs. The consistency score for data point $(q,r)$ is calculated as:
$\text{Cons}(q,r)=\frac{\sum_{i=1}^{X_{\text{MCQs}}}[R_{i}=r]}{X_{\text{MCQs}}}$ (4)
where $X_{\text{MCQs}}$ is the total number of MCQs, $R_{i}$ is model $M$’s response for the $i$-th MCQ, and $[R_{i}=r]$ yields 1 when the model’s response $R_{i}$ matches its original response $r$, and 0 otherwise. The consistency score $\text{Cons}(q,r)$ ranges from 0 to 1.
@@
measures a model’s consistency in its correct responses and is defined as:
@@
measures the consistency of an LLM when it provides wrong responses and is defined as:
@@
where $C_{\textit{wrong}}^{s}$/$C_{\textit{wrong}}^{u}$ refer to the consistency of an LLM when it provides wrong responses to questions about seen/unseen knowledge:
@@
where $r^{(w)}$ and $q^{(w)}$ denote wrong responses and their corresponding questions. $N_{wrong}$ and $L_{wrong}$ are the counts of wrong answers on $D_{\text{seen}}$ and $D_{\text{unseen}}$, respectively. $C_{\textit{wrong}}$ ranges from 0 to 1, and per Criterion 2.2, lower $C_{\textit{wrong}}$ is better.
@@ 2.3 Reliability (Factuality and Consistency)
Based on the criteria defined above, an LLM is reliable as a KB if it meets the following criteria when evaluated against factuality and consistency:
Criterion 3.1: For seen knowledge, an LLM should have a high rate of consistently correct responses and a low rate of consistently wrong responses. Criterion 3.2: For unseen knowledge, a LLM should have a high rate of uninformative or inconsistent responses.
We next quantify these criteria are with the following two metrics.
@@
NCCR ranges from $-1$ to $1$. NCCR values closer to 1 indicate an LLM is more reliable on seen knowledge. A negative NCCR suggests the model provides consistently wrong responses, while a positive NCCR suggests a preference for consistently correct responses.
@@
assesses whether an LLM is likely to provide uninformative or inconsistent wrong responses for questions about unseen knowledge. It is defined as: